What is the speed of a beam of pions if their average lifetime is measured to be 4.80 × 10-8 s? At rest, their lifetime is 2.00 × 10-8 s.
Answer #1

here by using the formula

Δt' = Δt / sqrt(1 - v^2/c^2)

4.80 * 10^-8 = 2 * 10^-8 / sqrt(1 - v^2/c^2)

sqrt(1 - v^2/c^2) = 0.4167

1 - v^2/c^2 = 0.1736

v^2/c^2 = 0.8264

v/c = 0.909

v = 0.909c

v = 0.909 * 3 * 10^8 = 2.727 * 10^8 m/s